[figure: 9.7 cm]

NEW MAGNETIC PHENOMENA IN VANADIUM HEXACYANOCROMATES

M. Lukáčová^{a}, L. F. Kiss^{b}, M. Maryško^{c},M. Mihalik^{a}, Z. Mitróová^{a}, J. Stopka^{d},A. Zentko^{a}, M. Zentková^{a}

aInstitute of Experimental Physics SAS, Watsonova 47, Košice, Slovak Republic

b Solid State Physics and Optics of HAS, Budapest, Hungary cInstitute of Physics ASCR, Na Slovance 2 Prague, Czech Republic dInstitute of Inorganic Chemistry ASCR, 250 68 Řež, Czech Republic

Magnetic properties of two vanadium hexacyanochromates belonging to the Prussian blue family of moleculebased magnetic materials are presented. It has been shown from magnetization measurements M(T) that the hexacyanocomplex 1 with V^{IV} orders ferrimagnetically below T_{c} = 120 K and the complex 2 with V^{III} orders ferrimagnetically below T_{c} = 180 K. Maximum in the magnetization at T_{max} = 8 K (sample 1) and at T_{max} = 37 K (sample 2) measured at H = 25 Oe shift to lower temperatures and becomes less pronounced at higher magnetic fields. The irreversibility in the field cooled and zerofieldcooled magnetization measurements performed up to magnetic field H = 1 kOe suggests a cluster glass behavior in the low temperature region for both samples. Both samples exhibit a strong paramagnetic resonance signal at g(1) = 1.84 and g(2) = 1.79 with the peak to peak linewidth ν(1) = 73 G and ν(2) = 944 G at room temperature. The results are discussed in relation with stoichiometry of three dimensional material and the nature of constituting metal ions.

[figure: 13.4 cm]

Subject category :

3. Transition Metals, Alloys and Compounds

Presentation mode : poster

Corresponding author : M. Lukáčová

Address for correspondence : Institute of Experimental Physics, Slovak Academy of Sciences, Watsonova 47, 043 53 Košice, Slovak Republic

Email address : broschm@saske.sk